# Operational Methods Applied to the Spherical Mean and X-Ray Transform

### v3

@@ 2 X-Ray Transform
In this short section, we apply the associated operator of the spherical mean to derive the inversion formula for the X-ray transform. This integral transform is very closely related to the Radon transform, and coincides with it in two dimensions. In general, the X-ray transform of a function $f$ is defined by integrating over lines in space; parameterizing the line as $L=\vec{x}+\ell\vec{\theta}$ for an initial point $\vec{x}$ and a direction $\vec{\theta}$ of the line, this can be written as
@@
and exchange the order of integration between the surface integral and the X-ray transform:
@@
The right integral is simply the spherical mean of $f(\vec{x})$, thus we can rewrite it with the operator (7) to get
@@
By rearranging this expression for $f(\vec{x})$, we find the following formal inversion formula:
@@
Of course, this formula depends on the interpretation of the operator $\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}$ and is of little use without specifying it. To make sense of the square root of the Laplacian, we multiply the numerator and denominator by $\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}$:
@@
where now the resulting operator, $1/\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}$, can be rigorously defined by its Fourier multiplier $\mathcal{F}\{1/\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}f(\vec{x})\}(\vec{k})\propto\lVert\vec{k}\rVert^{-1}\mathcal{F}\{f\}(\vec{k})$ and is known as the Riesz potential [21, 16]. It is given by the integral expression
$\frac{1}{\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}}f(\vec{x})=\frac{\Gamma(\frac{n-1}{2})}{2\pi^{\frac{n}{2}}\Gamma(\frac{n}{2})}\int_{\mathbb{R}^{n}}\mathrm{d}^{n}u\,\frac{f(\vec{u})}{\lVert\vec{x}-\vec{u}\rVert^{n-1}},$ (49)
which we can insert into our inversion formula to get the well-known expression [15]

### v4

@@ 2 X-Ray Transform
In this section, we apply the operator associated with the spherical mean to derive the inversion formula for the X-ray transform. This integral transform is closely related to the Radon transform and, in two dimensions, coincides with it. In general, we define the X-ray transform of a function $f$ by integrating $f$ along lines in space. Parameterizing a line by $L=\vec{x}+\ell\vec{\theta}$, where $\vec{x}$ denotes an initial point and $\vec{\theta}$ specifies the direction of the line, we can write
@@
and then exchange the order of integration between the surface integral and the X-ray transform:
@@
The inner integral on the right-hand side is simply the spherical mean of $f(\vec{x})$. We can therefore rewrite this expression using the operator form (7) as
@@
By rearranging this expression to solve for $f(\vec{x})$, we obtain the following formal inversion formula:
@@
This formula, however, depends on the interpretation of the operator $\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}$ and is therefore of limited practical use without further specification. To give meaning to the square root of the Laplacian, we multiply both the numerator and denominator by $\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}$, which yields
@@
Here, the resulting operator $1/\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}$ admits a rigorous definition through its Fourier multiplier, $\mathcal{F}\{1/\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}f(\vec{x})\}(\vec{k})\propto\lVert\vec{k}\rVert^{-1}\mathcal{F}\{f\}(\vec{k})$, and is known as the Riesz potential [21, 16]. It is given explicitly by the integral representation
$\frac{1}{\sqrt{-\mathop{{}\bigtriangleup}\nolimits\!_{\vec{x}}}}f(\vec{x})=\frac{\Gamma(\frac{n-1}{2})}{2\pi^{\frac{n}{2}}\Gamma(\frac{n}{2})}\int_{\mathbb{R}^{n}}\mathrm{d}^{n}u\,\frac{f(\vec{u})}{\lVert\vec{x}-\vec{u}\rVert^{n-1}}.$ (50)
Substituting this expression into the inversion formula, we recover the well-known result [15]